Ammonia (NH3) is an example of a Brønsted-Lowry Base.


                   i.     Define the Brønsted-Lowry acid-base theory.


                 ii.     What’s the pH of a solution of ammonia that has a concentration of 0.335 M? The Kb of ammonia is 1.8 × 10–5.


1
Expert's answer
2021-04-25T07:34:03-0400

1. Bronsted- Lowry acid base theory states that An acid is a proton donor while a base is a proton acceptor.

This implies that any compound that can transfer a proton to any compound is an ACID and a compound that can accept a proton is a BASE.


2.) The equation of reaction goes this: NH3+H20=> NH4+ + OH-

Initial concentration of NH3= 0.335M

Final concentration of NH3= 0.335-X

Concentration of NH4+ = X

Concentration of OH- = X


Kb = [NH4+][OH-]/[NH3]

1.8 x 10-5 = [X][X]/[0.335-X]


1.8 x 10-5 = X2/0.335-X

Assume that the value of X is infinitesimal.

1.8 x 10-5 = X2/0.335

X2= (1.8 x 10-5)(0.335)

X2= 6.03 x 10-6

Get the square root of both sides

X= 0.00246


[OH-] = X = 0.00246

pOH = -log[OH-]

pOH = -log[0.00246]

pOH= 2.61


pH+pOH= 14

pH = 14-pOH

pH= 14-2.61

pH = 11.39


Therefore the pH of the solution is 11.39